# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - AutoML from Basics to State-of-the-Art Techniques

### v2

@@ 19.2.1 VGG
VGG [96], proposed by the Visual Geometry Group at Oxford, is known for its simplicity and depth. The architecture consists of deep convolutional layers, with each layer followed by ReLU activation and max-pooling. What makes VGG special is that it uses small 3x3 filters throughout the network, combined with deep layers.
@@
6 self.conv1 = nn.Conv2d(3, 64, kernel_size=3, padding=1)
7 self.conv2 = nn.Conv2d(64, 128, kernel_size=3, padding=1)
8 self.fc = nn.Linear(128 * 56 * 56, 10) # Assuming input image is 224x224
@@ 19.2.2 Inception v1, v2, v3, v4
[…]
@@ 19.2.3 Xception
Xception [97] builds on the Inception architecture by replacing traditional convolutions with depthwise separable convolutions. This operation is more efficient because it first performs a spatial convolution for each channel individually and then combines them.
@@ 19.2.4 ResNet
ResNet [98], short for Residual Networks, introduces the concept of skip connections. These connections allow the model to learn residuals (the difference between the input and output of a layer), which solves the vanishing gradient problem, enabling very deep networks.
@@
1 class BasicBlock(nn.Module):
@@
4 self.conv1 = nn.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
5 self.conv2 = nn.Conv2d(out_channels, out_channels, kernel_size=3, padding=1)
6 self.skip_connection = nn.Sequential()
@@ 19.2.5 DenseNet
DenseNet [99] is another network that focuses on improving the gradient flow. In DenseNet, each layer receives inputs from all preceding layers, which enhances feature propagation and helps in alleviating the vanishing gradient problem.
@@
1 class DenseLayer(nn.Module):
@@
4 self.conv = nn.Conv2d(in_channels, growth_rate, kernel_size=3, padding=1)
@@ 20.1 Concept of Neural Architecture Search (NAS)
Neural Architecture Search (NAS) is an automated process used to design the structure of neural networks [100]. Traditional neural network design requires expert knowledge and time-consuming experimentation to determine the most effective architecture for a given task. NAS aims to automate this process, significantly reducing the effort required and often discovering architectures that outperform those designed manually [101].
NAS optimizes both the structure (topology) of a neural network and its hyperparameters. The search for the best architecture can be framed as an optimization problem, where the objective is to find the network configuration that maximizes performance on a given task, such as image classification or natural language processing [102].
@@ 20.2.3 Implementation and Applications of NASNet
5class NASNetCell(nn.Module):
@@
8 self.conv1 = nn.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
9 self.conv2 = nn.Conv2d(out_channels, out_channels, kernel_size=3, padding=1)
10 self.relu = nn.ReLU()
11 self.pool = nn.MaxPool2d(kernel_size=2, stride=2)
@@
18class NASNet(nn.Module):
@@
24 self.fc = nn.Linear(256 * 4 * 4, num_classes)
@@ 20.3 Other NAS Tools
  • DARTS (Differentiable Architecture Search): DARTS is a gradient-based NAS method that significantly reduces the computational cost of NAS [102]. Unlike traditional NAS methods, which require training many different architectures from scratch, DARTS allows for a continuous search space that can be optimized using gradients. This drastically reduces the number of required evaluations.
  • ENAS (Efficient Neural Architecture Search): ENAS is a reinforcement learning-based NAS method that focuses on efficiency [103]. It introduces the concept of a shared network, where different candidate architectures share weights during the training process. This reduces the computational overhead of NAS while still providing competitive results.
DARTS and ENAS represent two important trends in NAS research: improving the efficiency of the search process while maintaining high performance [100].
@@
4class DARTSCell(nn.Module):
@@
7 self.conv = nn.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
8 self.relu = nn.ReLU()
@@
14class DARTS(nn.Module):
@@
18 self.fc = nn.Linear(64 * 32 * 32, num_classes)
@@ 21.2 Auto-Keras
Auto-Keras [104] is an open-source AutoML library designed to make deep learning more accessible. It is built on top of Keras, but since this book focuses on PyTorch, we’ll discuss the concepts rather than the underlying TensorFlow framework.
@@ 22.1 Google Colab: Utilizing Free Resources
Google Colab [105] provides an excellent starting point for those without access to dedicated computational resources. It offers free access to CPU, GPU, and TPU environments for running machine learning models, including AutoML tasks.
@@ 22.3.2 Using SLURM to Schedule Jobs
SLURM [106] is another widely used workload manager for large compute clusters. Similar to PBS, you write a job script and submit it to the queue.
@@ 1. Scalability:
3from torch import nn, optim
@@
12class SimpleNN(nn.Module):
@@
15 self.fc1 = nn.Linear(4, 10)
16 self.fc2 = nn.Linear(10, 3)
@@ 23.2.1 Finance
11class FraudDetectionNN(nn.Module):
@@
14 self.fc1 = nn.Linear(10, 50)
15 self.fc2 = nn.Linear(50, 1)
@@ 23.3 Trends and Future Outlook
Another key trend is the integration of Explainable AI into AutoML workflows. In the future, AutoML systems will not only produce highly accurate models but also generate explanations that help stakeholders understand why a particular decision was made [107]. This will be particularly important in industries like finance and healthcare where trust and transparency are essential [108].
@@
As AutoML becomes more widespread, there will be a greater focus on ensuring that the models produced are ethical and fair [109]. There is already growing concern about bias in machine learning models, and AutoML will need to incorporate fairness constraints to ensure that the models it generates do not reinforce biases present in the data [110].

### v3

@@ 19.2.1 VGG
VGG [simonyan2015deepconvolutionalnetworkslargescale], proposed by the Visual Geometry Group at Oxford, is known for its simplicity and depth. The architecture consists of deep convolutional layers, with each layer followed by ReLU activation and max-pooling. What makes VGG special is that it uses small 3x3 filters throughout the network, combined with deep layers.
@@
6 self.conv1 = $n$.Conv2d(3, 64, kernel_size=3, padding=1)
7 self.conv2 = $n$.Conv2d(64, 128, kernel_size=3, padding=1)
8 self.fc = $n$.Linear(128 * 56 * 56, 10) # Assuming input image is 224x224
@@ 19.2.2 Inception v1, v2, v3, v4
Input Image1x1 Convolution3x3 Convolution5x5 ConvolutionMax-PoolingConcatenation
@@ 19.2.3 Xception
Xception [chollet2017xceptiondeeplearningdepthwise] builds on the Inception architecture by replacing traditional convolutions with depthwise separable convolutions. This operation is more efficient because it first performs a spatial convolution for each channel individually and then combines them.
@@ 19.2.4 ResNet
ResNet [he2015deepresiduallearningimage], short for Residual Networks, introduces the concept of skip connections. These connections allow the model to learn residuals (the difference between the input and output of a layer), which solves the vanishing gradient problem, enabling very deep networks.
@@
1 class BasicBlock($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
5 self.conv2 = $n$.Conv2d(out_channels, out_channels, kernel_size=3, padding=1)
6 self.skip_connection = $n$.Sequential()
@@ 19.2.5 DenseNet
DenseNet [huang2018denselyconnectedconvolutionalnetworks] is another network that focuses on improving the gradient flow. In DenseNet, each layer receives inputs from all preceding layers, which enhances feature propagation and helps in alleviating the vanishing gradient problem.
@@
1 class DenseLayer($n$.Module):
@@
4 self.conv = $n$.Conv2d(in_channels, growth_rate, kernel_size=3, padding=1)
@@ 20.1 Concept of Neural Architecture Search (NAS)
Neural Architecture Search (NAS) is an automated process used to design the structure of neural networks [elsken2019neural]. Traditional neural network design requires expert knowledge and time-consuming experimentation to determine the most effective architecture for a given task. NAS aims to automate this process, significantly reducing the effort required and often discovering architectures that outperform those designed manually [zoph2016neural].
NAS optimizes both the structure (topology) of a neural network and its hyperparameters. The search for the best architecture can be framed as an optimization problem, where the objective is to find the network configuration that maximizes performance on a given task, such as image classification or natural language processing [liu2018darts].
@@ 20.2.3 Implementation and Applications of NASNet
5class NASNetCell($n$.Module):
@@
8 self.conv1 = $n$.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
9 self.conv2 = $n$.Conv2d(out_channels, out_channels, kernel_size=3, padding=1)
10 self.relu = $n$.ReLU()
11 self.pool = $n$.MaxPool2d(kernel_size=2, stride=2)
@@
18class NASNet($n$.Module):
@@
24 self.fc = $n$.Linear(256 * 4 * 4, num_classes)
@@ 20.3 Other NAS Tools
  • DARTS (Differentiable Architecture Search): DARTS is a gradient-based NAS method that significantly reduces the computational cost of NAS [liu2018darts]. Unlike traditional NAS methods, which require training many different architectures from scratch, DARTS allows for a continuous search space that can be optimized using gradients. This drastically reduces the number of required evaluations.
  • ENAS (Efficient Neural Architecture Search): ENAS is a reinforcement learning-based NAS method that focuses on efficiency [pham2018efficient]. It introduces the concept of a shared network, where different candidate architectures share weights during the training process. This reduces the computational overhead of NAS while still providing competitive results.
DARTS and ENAS represent two important trends in NAS research: improving the efficiency of the search process while maintaining high performance [elsken2019neural].
@@
4class DARTSCell($n$.Module):
@@
7 self.conv = $n$.Conv2d(in_channels, out_channels, kernel_size=3, padding=1)
8 self.relu = $n$.ReLU()
@@
14class DARTS($n$.Module):
@@
18 self.fc = $n$.Linear(64 * 32 * 32, num_classes)
@@ 21.2 Auto-Keras
Auto-Keras [jin2019autokerasefficientneuralarchitecture] is an open-source AutoML library designed to make deep learning more accessible. It is built on top of Keras, but since this book focuses on PyTorch, we’ll discuss the concepts rather than the underlying TensorFlow framework.
@@ 22.1 Google Colab: Utilizing Free Resources
Google Colab [googlecolab] provides an excellent starting point for those without access to dedicated computational resources. It offers free access to CPU, GPU, and TPU environments for running machine learning models, including AutoML tasks.
@@ 22.3.2 Using SLURM to Schedule Jobs
SLURM [yoo2003slurm] is another widely used workload manager for large compute clusters. Similar to PBS, you write a job script and submit it to the queue.
@@ 1. Scalability:
3from torch import $n$, optim
@@
12class SimpleNN($n$.Module):
@@
15 self.fc1 = $n$.Linear(4, 10)
16 self.fc2 = $n$.Linear(10, 3)
@@ 23.2.1 Finance
11class FraudDetectionNN($n$.Module):
@@
14 self.fc1 = $n$.Linear(10, 50)
15 self.fc2 = $n$.Linear(50, 1)
@@ 23.3 Trends and Future Outlook
Another key trend is the integration of Explainable AI into AutoML workflows. In the future, AutoML systems will not only produce highly accurate models but also generate explanations that help stakeholders understand why a particular decision was made [drozdal2020trust]. This will be particularly important in industries like finance and healthcare where trust and transparency are essential [arrieta2020explainable].
@@
As AutoML becomes more widespread, there will be a greater focus on ensuring that the models produced are ethical and fair [hazan2020autml]. There is already growing concern about bias in machine learning models, and AutoML will need to incorporate fairness constraints to ensure that the models it generates do not reinforce biases present in the data [mehrabi2021survey].